1. Rights and privileges of common stockholders Larry Nelson holds 1,000 shares of General Electric (GE) common stock. As a sTrue False $60,000 $100,000 Larry also holds 2,000 shares of common sto any that only has 20,000 shares outstanding. The compLarry also holds 2,000 sh $50.00 per share. The co $40.00 per share. Larry $120,000 mon stock in a company that only has 20,0Larry also holds 2,000 shares of common stock in a company that only has 20,000 shares outstanding. The companys stock curreLarry also holds 2,000 shares of common stock in a company that only has 20,000 shares outstanding. The companys stock curreLarry also holds 2,000 shares of common stock in a company that only has 20,000 shares outstanding. The companys stock curre
Answer #1

1) False , Larry have common stocks , so only after preferred shareholders are paid dividend , the common stockholders could be paid so they won't receive it together

2)$100,000 , as stock price is $50 and he has 2000 shares , so his investment is worth $100,000

3) $96000, the weighted average of the stock price is done as shown below
w1= 20000 , p1= $50 , w2=5000 , p2 = $40

Now Weighted Average Price = (w1p1 + w2p2)/(w1+w2)

= (20000*50 + 5000*40)/ 25000

=$48
He owned total 2000 shares , so his investment is worth $48* 2000= $96000

4) Dilution , the scenario is an example of dilution in which company issues more equity shares and dilute existing shareholders percentage .

5) preemptive rights, this are rights through which existing shareholders could maintain their percentage ownership of company by buying a proportionate number of shares in the future

6) $120,000, We had exercised this provision he would have 2500 shares(10% of total 25000) and share prices would $48 , so total would be 48*2500 = $120,000